# Data-Driven Global Sensitivity Analysis for Engineering Design Based on Individual Conditional Expectations

### v1

@@ Abstract
Explainable machine learning techniques have gained increasing attention in engineering applications, especially in aerospace design and analysis, where understanding how input variables influence data-driven models is essential. Partial Dependence Plots (PDPs) are widely used for interpreting black-box models by showing the average effect of an input variable on the prediction. However, their global sensitivity metric can be misleading when strong interactions are present, as averaging tends to obscure interaction effects. To address this limitation, we propose a global sensitivity metric based on Individual Conditional Expectation (ICE) curves. The method computes the expected feature importance across ICE curves, along with their standard deviation, to more effectively capture the influence of interactions. We provide a mathematical proof demonstrating that the PDP-based sensitivity is a lower bound of the proposed ICE-based metric under truncated orthogonal polynomial expansion. In addition, we introduce an ICE-based correlation value to quantify how interactions modify the relationship between inputs and the output. Comparative evaluations were performed on three cases: a 5-variable analytical function, a 5-variable wind-turbine fatigue problem, and a 9-variable airfoil aerodynamics case, where ICE-based sensitivity was benchmarked against PDP, SHapley Additive exPlanations (SHAP), and Sobol’ indices. The results show that ICE-based feature importance provides richer insights than the traditional PDP-based approach, while visual interpretations from PDP, ICE, and SHAP complement one another by offering multiple perspectives.
@@ 1 Introduction
High-fidelity simulations are essential in modern engineering design. Therefore, aerospace engineers often rely on computational fluid dynamics to assess and optimize aerodynamic performance, enabling accurate prediction of behavior under varied conditions and supporting informed, performance- and safety-driven decisions [1]. As engineering problems grow more complex, especially in aerospace, precise simulation data is important. Yet high-fidelity simulations remain challenging for design exploration because engineers can only afford to run a limited number of computationally intensive simulations. To address this, a surrogate model that estimates the relationship between input or design variables with the Quantity of Interest (QoI) plays an important role in accomplishing this task. The surrogate can be used for various purposes such as optimization [2] and uncertainty analysis [3]. Various types of surrogate models exist, among popular ones in engineering, including polynomial regression (including polynomial chaos expansion (PCE) [4, 5]), Kriging/Gaussian process regression [6, 7], and deep neural network [8] to reduce the computational cost of inference.
@@
Sensitivity analysis, particularly Global Sensitivity Analysis (GSA), supports engineering endeavours by identifying which variables are most and least influential [9]. In the context of engineering design exploration, this information is important for setting design priorities, such as determining which variables should be prioritized. For instance, in the field of aerodynamics, GSA aids in analyzing the impact of geometric or kinematic parameters on the aerodynamic performance of a multi-element airfoil [10], a hovering flapping wing [11], and a high-pressure capturing wing [12]. Common techniques in engineering include variance-based decomposition methods, such as Sobol' indices [13], which measure the individual effects of input variables and their interactions on the QoI. Although variance-based decomposition is valuable, it only quantifies how much each input contributes to output variance but does not directly reveal the input–output functional form, such as local nonlinearity or the detailed structure of interactions. This is why we argue that visual tools such as PDPs and ICE curves provide complementary information to variance decomposition.
@@
Besides GSA, sometimes it is also of interest to understand the behavior of the black-box function when the input of interest is changed. In artificial intelligence and Machine Learning (ML), explainability serves several purposes, including building trust in the model and enhancing transparency, which is especially crucial in fields like medicine [14], and knowledge discovery [15]. Even in engineering, transparency is also vital, as it allows engineers and researchers to verify that the ML model is consistent with established domain knowledge. In this paper, we aim to utilize explainability methods for knowledge discovery in engineering design, viewing the ML model not just as a predictive tool, but also as a source of valuable insights. Recently, there has been growing interest in unifying the frameworks of sensitivity analysis and explainable ML [16].
@@
Techniques from explainable ML and graphical statistics such as Partial Dependence Plot (PDP) [17], Individual Conditional Expectation (ICE) [18], and SHapley Additive exPlanations (SHAP) [19] can be leveraged to understand the behavior of the black-box function. These methods provide insights into how input variables influence the predictive model and interact with each other. Such information is invaluable in engineering and science, helping engineers and scientists extract meaningful insights from the constructed model. The applications of explainable machine learning in engineering are now gaining significant attention, with examples not only in aerospace and mechanical engineering [20, 21, 22, 23], but also in material science [24] and civil engineering [25].
@@ 2 Motivations
The widespread use of high-fidelity simulations in aerospace means that engineers must do more than predict the QoI: they must quantify uncertainty sources and interpret black-box surrogates. This section briefly surveys the convergence of GSA and XAI, and exposes that engineering requires sensitivity analysis that goes beyond classical variance decomposition.
@@ 2.1 Global Sensitivity Analysis in Engineering Design
In high-dimensional engineering problems, GSA is essential for knowledge discovery and variable screening; the dominant framework is variance-based sensitivity methods, and are widely considered the “gold standard” in various fields [13, 30]. For high-dimensional problems, estimating variance contributions can be prohibitive in terms of computational time, especially for dummy Monte Carlo implementation [31]. This has led to the popularity of Derivative-based Global Sensitivity Measures (DGSM) [32], which use gradients to bound sensitivity indices, and Leave-one-covariate-out metrics [33], which focus on predictive power rather than variance. Still, many different techniques may alleviate the problem of sensitivity indices computation in high dimensions [34, 35], but with a small number of data points to learn on in high dimensions, these indices remain difficult to estimate [36].
Concerning the computational cost, it is worth noting that many authors have proposed leveraging surrogate models, such as PCE, to perform GSA for complex systems in a tractable and time-efficient manner [30]. Still, when GSA is performed on data-driven surrogates, sensitivity estimates generally mix up input variability and surrogate approximation error unless surrogate uncertainty is explicitly modeled and propagated [30]. Even with many advances, sensitivity indices only provide a single scalar measure of influence, indicating that a variable is important but failing to describe the mechanism of influence. For instance, an index will not reveal how the output behaves across the input’s range (e.g., if the response is monotonically increasing, quadratic, or exhibits a threshold). Understanding this detailed shape of the output response is a prerequisite for making informed design trade-offs [37].
@@ 2.2 The Rise of Explainable AI for Engineering
Explainable Artificial Intelligence (XAI) can complement scalar GSA by exposing the functional and instance-level behaviour of complex surrogates used in aerospace design. In practice, XAI delivers two complementary views: local explanations (why a specific prediction was made for a specific input) and global explanations (how the model behaves across the input space). In engineering workflows, local explanations support root-cause analysis and failure investigation, while global views guide design exploration and trade-offs. The most used technique in such a context is SHAP [19] that provides principled local attributions based on cooperative game theory. Importantly, SHAP (local) is distinct from Shapley effects (a global GSA measure): aggregating per-instance SHAP values (e.g., mean absolute SHAP) yields a heuristic global ranking but can mask heterogeneity across the operational design space.
PDP [17] estimates a variable’s marginal profile by averaging the model output over other inputs. While PDP can reveal the shape of a feature’s effect (e.g., linear, nonlinear), it does not directly equate to first-order variance decomposition (such as Sobol’ main effects), because it averages across both main and interaction contributions rather than isolating just the variance attributed to that feature [38]. PDPs are intuitive and widely used to visualise input-output trends, but their averaging hides heterogeneous interactions: when an input has opposite effects in different regions, PDPs can cancel these effects and produce a misleadingly flat profile, a serious hazard in safety-critical cases.
ICE curves [18] plot per-instance response trajectories, revealing heterogeneity that PDP averages obscure. ICE therefore offers richer diagnostic power for aerospace problems, in which understanding is critical. However, their utility in automated engineering design pipelines is hindered by the absence of standardised quantitative metrics or scalar summaries necessary for efficient variable ranking. A key research priority for broader engineering adoption is the development of robust ICE-based scalar metrics, which is the topic of the present paper.
These XAI tools, when used with awareness of their assumptions and computational costs, form a complementary toolkit for interpretable aerospace design.
@@ 2.3 Positioning of the Present Work
To address the limitations of the “cancellation effect” in PDPs, Goldstein et al. [18] introduced ICE plots. Instead of plotting the average, ICE plots display the functional relationship for every instance in the dataset, resulting in a set of curves that reveal the full dispersion of model behavior. There is currently a lack of formalised metrics that translate the visual richness of ICE curves into scalar sensitivity indices suitable for automated ranking in engineering design. To that end, this work bridges the gap between the rigorous quantification of variance-based GSA and the visual interpretability of ICE. We propose a novel ICE-based sensitivity metric that quantifies the dispersion of ICE curves. By calculating the standard deviation of the conditional expectations, we derive a metric that:
@@
    Captures interaction effects that are mathematically canceled out in standard PDP-based metrics.
@@
    Provides a theoretical lower bound to ICE-based sensitivity.
@@
This approach allows engineers to retain the intuitive visualization of response curves while ensuring that critical interactions are not masked by averaging artifacts.
3 Background and previous works
@@ 3.1 Surrogate model
Let $\boldsymbol{x}=(x_{1},x_{2},\ldots,x_{m})^{T}$ be the vector of the input of interest, where $m$ is the number of input variables. Assuming independence between input variables, let us also define the domain of the input variables as $\boldsymbol{\Omega}=\prod_{i=1}^{m}\Omega_{j}$, where $\Omega_{j}$ is the domain for the $i$-th input variable. Let us also define the joint density function $\rho(\boldsymbol{x})=\prod_{j=1}^{m}\rho_{x_{j}}(x_{j})$, where $\rho_{x_{j}}$ where $j=1,\ldots,m$ represents the marginal density. Also, let $y=f(\boldsymbol{x})$ be a black-box function that relates $\boldsymbol{x}$ to the QoI $y$. A surrogate model $\hat{f}(\boldsymbol{x})$ essentially approximates the relationship between $y$ and $\boldsymbol{x}$ using an estimation $\hat{f}(\boldsymbol{x})$, based on the given data set $\mathcal{D}=\{(\boldsymbol{x}^{(i)},y^{(i)})\}_{i=1}^{n}$, where $n$ is the size of the data set. In this paper, we use the data-driven non-intrusive PCE as our surrogate model of choice, although other models could also be applicable.
The PCE model is briefly outlined below, with more detailed explanations available in dedicated literature (e.g.[39]). The expansion reads as
@@
where the $\boldsymbol{\Psi}_{\boldsymbol{\gamma}}$’s are multidimensional polynomials that are orthonormal with respect to the joint density $\rho(x)$, and $\boldsymbol{\gamma}=\{\gamma_{1},\ldots,\gamma_{m}\}$ with $\gamma_{i}\geq 0$ for each $i$, denotes the individual indices which are members of $\mathbb{A}_{p}$. The coefficients $\boldsymbol{\beta}=\{\beta_{\gamma}\}$ corresponds to the terms of the PCE coefficients, where $P$ is the cardinality of $\mathbb{A}_{p}$.
The index set $\mathbb{A}_{p}$ is generated via the hyperbolic truncation as follows [39]:
$\mathbb{A}_{p}\equiv\{\boldsymbol{\gamma}\in\mathbb{N}^{m}:\|\boldsymbol{\gamma}\|_{q}\leq p\},$ (2)
where
$\|\boldsymbol{\gamma}\|_{q}=\bigg(\sum_{i=1}^{m}\gamma_{i}^{q}\bigg)^{1/q},0<q\leq 1,$ (3)
where $p$ is the polynomial order and $q$ controls the sparsity of the expansion. In this paper, we use $q=1$ since we deal with inputs of medium dimensionality.
@@
In this study, we employ the data-driven PCE module from UQLab [40], whereby the orthogonal polynomials for the engineering problems are constructed directly from the available data by inferring the probability distributions first.
@@ 3.2 Partial Dependence Plots and Individual Conditional Expectations
Let us define $1:m]=\{1,2,\ldots,m\}$. Next, let $S\subset[1:m]$ represent a subset of these indices, and let $C=[1:m]\setminus S$ denote the complementary set of indices not included in $S$. Also, let $\boldsymbol{x}_{S}=(x_{i}|i\in S)$ and $\boldsymbol{x}_{C}=(x_{i}|i\in C)$ represent the vectors of inputs corresponding to $S$ and $C$, respectively. Specifically, we want to build the partial dependence function of the output of the variables indexed by $S$. Let $\Omega_{S}$ be the subdomain for the variables indexed by $S$, i.e., $\boldsymbol{\Omega}_{S}=\prod_{i\in S}\Omega_{i}$. Similarly, $\boldsymbol{\Omega}_{C}=\prod_{i\in C}\Omega_{i}$. Finally, we define the joint probability density functions for these subsets as $\boldsymbol{\rho}_{\boldsymbol{x}_{S}}=\prod_{i\in S}\rho_{x_{i}}(x_{i})$ and $\boldsymbol{\rho}_{\boldsymbol{x}_{C}}=\prod_{i\in C}\rho_{x_{i}}(x_{i})$.
@@
$\hat{f}_{pdp}(\boldsymbol{x}_{S})=\mathbb{E}_{\boldsymbol{x}_{C}}\left[\hat{f}\left(\boldsymbol{x}_{S},\boldsymbol{x}_{C}\right)\right]=\int_{\boldsymbol{\Omega}_{C}}\hat{f}(\boldsymbol{x}_{S},\boldsymbol{x}_{C})\boldsymbol{\rho}_{\boldsymbol{x}_{C}}(\boldsymbol{x}_{C})d\boldsymbol{x}_{C},$ (4)
@@
$\hat{f}_{pdp}(\boldsymbol{x}_{S})\approx\frac{1}{n_{mc}}\sum_{i=1}^{n_{mc}}\hat{f}(\boldsymbol{x}_{S},\boldsymbol{x}_{C}^{(i)}),$ (5)
where $n_{mc}$ is the size of the random sampling set and $\boldsymbol{x}_{C}^{(i)}$ is the $i$-th realization of samples drawn from $\boldsymbol{\rho}_{\boldsymbol{x}_{C}}(\boldsymbol{x}_{C})$.
Because PDP averages over other features, it can hide interaction effects; For example, showing a flat line when equal portions of data have positive and negative correlations with the output. This issue is resolved by displaying the ICE curves together with the main PDP [18]. A single ICE curve essentially represents a single $\hat{f}(\boldsymbol{x}_{S},\boldsymbol{x}_{C}^{(i)})$ (see Eq. (5)). That is, a single ICE plot illustrates the effect of changing the input of interest while keeping the other variables constant. To model interaction is to capture non-additivity and more complex effects. Visually, ICE diagnostics reveal the structure of interactions: heterogeneity in curve slopes indicates that the effect of $x_{j}$ depends on $\boldsymbol{x_{C}}$.
Anchoring the PDP and ICE curves at a specific point can better visualize the interactions. To that end, in depicting PDP and ICE curves, we anchor the curves at the centre or mean of the input space. The anchored partial dependence function (i.e., $\hat{f}_{pdp}^{\text{anc}}(\boldsymbol{x}_{S},\boldsymbol{x_{C}})$) is reads as
$\hat{f}_{pdp}^{\text{anc}}(\boldsymbol{x}_{S},\boldsymbol{x_{C}})=\hat{f}_{pdp}(\boldsymbol{x}_{S})-\boldsymbol{1}\hat{f}_{pdp}(\boldsymbol{x}_{S}^{a},\boldsymbol{x}_{C}),$ (6)
where $\boldsymbol{x}_{S}^{a}=\mathbb{E}(\boldsymbol{x}_{S})$ is the anchor point. The ICE curves are also anchored using the same principle. In the discussions and results that follow, we always depict the PDP and ICE curves anchored through this procedure.

### v3

# Global Sensitivity Analysis for Engineering Design Based on Individual Conditional Expectations

@@ Abstract
Explainable machine learning techniques have gained increasing attention in engineering applications, especially in aerospace design and analysis, where understanding how input variables influence predictive models is essential. Partial Dependence Plots (PDPs) are widely used for interpreting black-box models by showing the average effect of an input variable on the prediction. However, their global sensitivity metric can be misleading when strong interactions are present, as averaging tends to obscure interaction effects. To address this limitation, we propose a global sensitivity metric based on Individual Conditional Expectation (ICE) curves. The method computes the expected feature importance across ICE curves, along with their standard deviation, to more effectively capture the influence of interactions. The proposed metrics are model-agnostic and can be applied to any predictive model, including but not limited to surrogate models. Furthermore, we provide a mathematical proof demonstrating that the PDP-based sensitivity is a lower bound of the proposed ICE-based metric under additive and multiplicative separability. In addition, we introduce an ICE-based correlation value to quantify how interactions modify the relationship between inputs and the output. Comparative evaluations were performed on three cases: a 5-variable analytical function, a 5-variable wind-turbine fatigue problem, and a 9-variable airfoil aerodynamics case, where ICE-based sensitivity was benchmarked against PDP, SHapley Additive exPlanations (SHAP), and Sobol’ indices. The results show that ICE-based feature importance provides richer insights than the traditional PDP-based approach, while visual interpretations from PDP, ICE, and SHAP complement one another by offering multiple perspectives.
@@ 1 Introduction
High-fidelity simulations are essential in modern engineering design. Therefore, aerospace engineers often rely on computer simulations to assess and optimize performance, enabling accurate prediction of behavior under varied conditions and supporting informed, performance- and safety-driven decisions [47]. Yet high-fidelity simulations remain challenging for design exploration because engineers can only afford to run a limited number of computationally intensive simulations. To address this, a surrogate model that estimates the relationship between input or design variables with the Quantity of Interest (QoI) plays an important role in accomplishing this task. The surrogate can be used for various purposes such as optimization [4] and uncertainty analysis [9]. Various types of surrogate models exist, among popular ones in engineering, including polynomial regression (including Polynomial Chaos Expansion (PCE) [16]), Kriging/Gaussian process regression [39, 46], and deep neural network [30] to reduce the computational cost of inference.
@@
Sensitivity analysis, particularly Global Sensitivity Analysis (GSA), supports engineering endeavours by identifying which variables are most and least influential [41]. In the context of engineering design exploration, this information is important for setting design priorities, such as determining which variables should be prioritized. For instance, in the field of aerodynamics, GSA aids in analyzing the impact of geometric or kinematic parameters on the aerodynamic performance of a multi-element airfoil [38], a hovering flapping wing [29], and a high-pressure capturing wing [55]. Common techniques in engineering include variance-based decomposition methods, such as Sobol’ indices [51], which measure the individual effects of input variables and their interactions on the QoI. Although variance-based decomposition is valuable, it only quantifies how much each input contributes to output variance but does not directly reveal the input–output functional form, such as local nonlinearity or the detailed structure of interactions.
@@
Besides GSA, sometimes it is also of interest to understand the behavior of the black-box function when the input of interest is changed. In artificial intelligence and Machine Learning (ML), explainability serves several purposes, including building trust in the model and enhancing transparency [3, 43]. Even in engineering, transparency is also vital, as it allows engineers and researchers to verify that the ML model is consistent with established domain knowledge. In this paper, we aim to utilize explainability methods for knowledge discovery in engineering design, viewing the ML model not just as a predictive tool, but also as a source of valuable insights. Recently, there has been growing interest in unifying the frameworks of sensitivity analysis and explainable ML [48].
@@
Techniques from explainable ML and graphical statistics such as Partial Dependence Plot (PDP) [14], Individual Conditional Expectation (ICE) [17], and SHapley Additive exPlanations (SHAP) [33] can be leveraged to understand the behavior of the black-box function. These methods provide insights into how input variables influence the predictive model and interact with each other. Such information is invaluable in engineering and science, helping engineers and scientists extract meaningful insights from the constructed model. The applications of explainable ML in engineering are now gaining significant attention, with examples not only in aerospace and mechanical engineering [50, 49, 54, 26], but also in other fields such as material science [37].
@@ 2 Motivations
The widespread use of high-fidelity simulations in aerospace means that engineers must do more than predict the QoI: careful interpretation of black-box surrogate models is now required. This section briefly surveys the convergence of GSA and Explainable Artificial Intelligence (XAI), and exposes that engineering requires sensitivity analysis that goes beyond classical variance decomposition.
@@ 2.1 Global Sensitivity Analysis in Engineering Design
In high-dimensional engineering problems, GSA is essential for knowledge discovery and variable screening; the dominant framework is variance-based sensitivity methods that are widely considered the “gold standard” in various fields [51, 53]. For high-dimensional problems, estimating variance contributions can be prohibitive in terms of computational time, especially for dummy Monte Carlo implementation [12]. This has led to the popularity of Derivative-based Global Sensitivity Measures (DGSM) [24], which use gradients to bound sensitivity indices, and Leave-one-covariate-out metrics [31], which focus on predictive power rather than variance. Still, many different techniques may alleviate the problem of sensitivity indices computation in high dimensions [32, 52], but with a small number of data points to learn on in high dimensions, these indices remain difficult to estimate [2].
Concerning the computational cost, it is worth noting that many authors have proposed leveraging surrogate models, such as PCE, to perform GSA for complex systems in a tractable and time-efficient manner [53]. Even with many advances, sensitivity indices only provide a single scalar measure of influence, indicating that a variable is important but failing to describe the mechanism of influence. For instance, if the response is monotonically increasing, quadratic, or exhibits a threshold, an index will not reveal how the output behaves across the input’s range. Understanding this detailed shape of the output response is a prerequisite for making informed design trade-offs [7].
@@ 2.2 The Rise of Explainable AI for Engineering
XAI can complement scalar GSA by exposing the functional and instance-level behavior of complex surrogates used in aerospace design. In practice, XAI delivers two complementary views: local explanations (why a specific prediction was made for a specific input) and global explanations (how the model behaves across the input space). In engineering workflows, local explanations support root-cause analysis and failure investigation, while global views guide design exploration and trade-offs. The most used technique in such a context is SHAP [33] that provides principled local attributions based on cooperative game theory.
On the other hand, PDP [14] estimates a variable’s marginal profile by averaging the model output over other inputs. PDPs are intuitive and widely used to visualise input-output trends, but their averaging hides heterogeneous interactions: when an input has opposite effects in different regions, PDPs can cancel these effects and produce a misleadingly flat profile, a serious hazard in safety-critical cases. On the other hand, ICE curves [17] plot per-instance response trajectories, revealing heterogeneity that PDP averages obscure. ICE therefore offers richer diagnostic power for engineering problems, in which understanding is critical. However, their utility in automated engineering design pipelines is hindered by the absence of standardised quantitative metrics or scalar summaries necessary for efficient variable ranking. A key research priority for broader engineering adoption is the development of ICE-based scalar metrics, which is the topic of the present paper. These XAI tools, when used with awareness of their assumptions and computational costs, form a complementary toolkit for interpretable aerospace design.
@@ 2.3 Positioning of the Present Work
There is currently a lack of formalised metrics that translate the visual richness of ICE curves into scalar sensitivity indices suitable for automated ranking in engineering design. To that end, this work bridges the gap between GSA and the visual interpretability of ICE. We propose novel ICE-based sensitivity metrics that are based on the dispersion of ICE curves. By focusing on conditional expectations, we derive metrics that:
@@
    Capture interaction effects that are mathematically cancelled out in standard PDP-based metrics.
@@
    Provides a theoretical lower bound to the proposed ICE-based sensitivity metric.
@@
This approach allows engineers to retain the intuitive visualization of response curves while ensuring that critical interactions are not masked by averaging artefacts. This paper introduces sensitivity analysis tools derived from local conditional behavior that bridge the computational gap between local and global sensitivity analysis and provide a quantitative complement to graphical methods. Also, the methods proposed in this paper apply to any response function, including black-box models and surrogate representations. For computationally expensive models, such as those arising from computational fluid dynamics simulations, the proposed GSA metrics can be evaluated using a surrogate model, provided that the surrogate achieves sufficient predictive accuracy to ensure reliable sensitivity estimates and knowledge extraction.
3 Surrogate Modeling for Global Sensitivity and Explainability Analysis
@@ 3.1 Surrogate model
Let $\boldsymbol{x}=(x_{1},x_{2},\ldots,x_{m})^{T}$ be the vector of the input of interest, where $m$ is the number of input variables. Assuming independence between input variables, let us define the domain of the input variables as $\boldsymbol{\Omega}=\prod_{j=1}^{m}\Omega_{j}$, where $\Omega_{j}$ is the domain for the $j$-th input variable. Let us also define the joint density function $\boldsymbol{\pi}(\boldsymbol{x})=\prod_{j=1}^{m}\pi_{x_{j}}(x_{j})$, where $\pi_{x_{j}}$ represents the marginal density of $x_{j}$. Also, let $y=f(\boldsymbol{x})$ be a black-box function that relates $\boldsymbol{x}$ to the QoI $y$. A surrogate model $\hat{f}(\boldsymbol{x})$ essentially approximates the relationship between $y$ and $\boldsymbol{x}$, based on the given data set $\mathcal{D}=\{(\boldsymbol{x}^{(i)},y^{(i)})\}_{i=1}^{n}$, where $n$ is the size of the data set. In this paper, we use the data-driven non-intrusive PCE as our surrogate model of choice, although other models could also be applicable.
While more detailed explanations are given in [5], the PCE model is briefly outlined below. The expansion reads as
@@
where the $\boldsymbol{\Psi}_{\boldsymbol{\gamma}}$’s are multidimensional polynomials that are orthonormal with respect to the joint density $\boldsymbol{\pi}(\boldsymbol{x})$, and $\boldsymbol{\gamma}=\{\gamma_{1},\ldots,\gamma_{m}\}$ with $\gamma_{i}\geq 0$ for each $i$, denotes the individual indices which are members of $\mathbb{A}_{p}$. The coefficients $\boldsymbol{\beta}=\{\beta_{\gamma}\}$ corresponds to the terms of the PCE coefficients, where $P$ is the cardinality of $\mathbb{A}_{p}$. The index set $\mathbb{A}_{p}$ is constructed using a total-order truncation [5]:
$\mathbb{A}_{p}\equiv\{\boldsymbol{\gamma}\in\mathbb{N}^{m}:\|\boldsymbol{\gamma}\|_{1}\leq p\},$ (2)
where $p$ denotes the polynomial order and
$\|\boldsymbol{\gamma}\|_{1}=\sum_{i=1}^{m}\gamma_{i}.$ (3)
@@
The method and the framework proposed in this paper, explained in detail in Sect. 4, are surrogate-agnostic and do not rely on any specific properties of PCE. PCE is adopted here due to its spectral and additive structure, which facilitates the theoretical derivation of the inequality in Eq. (19) and enables analytical tractability, as detailed in Appendix A. In addition, for the engineering demonstrations considered in this paper, the PCE surrogate achieves excellent predictive accuracy. In this study, we employ the data-driven PCE module from UQLab [35], whereby the orthogonal polynomials for the engineering problems are constructed directly from the available data by inferring the probability distributions first.
@@ 3.2 Partial Dependence Plots and Individual Conditional Expectations
Let us define $1:m]=\{1,2,\ldots,m\}$. Next, let $S\subset[1:m]$ represent a subset of these indices, and let $C=[1:m]\setminus S$ denote the complementary set of indices not included in $S$. Also, let $\boldsymbol{x}_{S}=(x_{j}|j\in S)$ and $\boldsymbol{x}_{C}=(x_{j}|j\in C)$ represent the vectors of inputs corresponding to $S$ and $C$, respectively. Specifically, we want to build the partial dependence function of the output of the variables indexed by $S$. Let $\Omega_{S}$ be the subdomain for the variables indexed by $S$, i.e., $\boldsymbol{\Omega}_{S}=\prod_{j\in S}\Omega_{j}$. Similarly, $\boldsymbol{\Omega}_{C}=\prod_{j\in C}\Omega_{j}$. Finally, we define the joint probability density functions for these subsets as $\boldsymbol{\pi}_{\boldsymbol{x}_{S}}=\prod_{j\in S}\pi_{x_{j}}(x_{j})$ and $\boldsymbol{\pi}_{\boldsymbol{x}_{C}}=\prod_{j\in C}\pi_{x_{j}}(x_{j})$.
@@
$\hat{f}_{pdp}(\boldsymbol{x}_{S})=\mathbb{E}_{\boldsymbol{x}_{C}}\left[\hat{f}\left(\boldsymbol{x}_{S},\boldsymbol{x}_{C}\right)\right]=\int_{\boldsymbol{\Omega}_{C}}\hat{f}(\boldsymbol{x}_{S},\boldsymbol{x}_{C})\boldsymbol{\pi}_{\boldsymbol{x}_{C}}(\boldsymbol{x}_{C})d\boldsymbol{x}_{C},$ (4)
@@
$\hat{f}_{pdp}(\boldsymbol{x}_{S})\approx\frac{1}{N}\sum_{i=1}^{N}\hat{f}(\boldsymbol{x}_{S},\boldsymbol{x}_{C}^{(i)}),$ (5)
where $N$ is the size of the random sampling set and $\boldsymbol{x}_{C}^{(i)}$ is the $i$-th realization of samples drawn from $\boldsymbol{\pi}_{\boldsymbol{x}_{C}}(\boldsymbol{x}_{C})$.
Because PDP averages over other features, it can hide interaction effects; For example, showing a flat line when equal portions of data have positive and negative correlations with the output. This issue is resolved by displaying the ICE curves together with the main PDP [17]. A single ICE curve essentially represents a single $\hat{f}(\boldsymbol{x}_{S},\boldsymbol{x}_{C}^{(i)})$ (see Eq. (5)). That is, a single ICE plot illustrates the effect of changing the input of interest while keeping the other variables constant. To model interaction is to capture non-additivity and more complex effects. Visually, ICE diagnostics reveal the structure of interactions: heterogeneity in curve slopes indicates that the effect of $\boldsymbol{x_{S}}$ depends on $\boldsymbol{x_{C}}$.
Anchoring the PDP and ICE curves at a specific point can better visualize the interactions. In depicting PDP and ICE curves, we anchor the curves at the centre or mean of the input space. The anchored partial dependence function (i.e., $\hat{f}_{pdp}^{\text{anc}}(\boldsymbol{x}_{S}$) is reads as
$\hat{f}_{pdp}^{\text{anc}}(\boldsymbol{x}_{S})=\hat{f}_{pdp}(\boldsymbol{x}_{S})-\hat{f}_{pdp}(\boldsymbol{x}_{S}^{a}),$ (6)
where $\boldsymbol{x}_{S}^{a}=\mathbb{E}(\boldsymbol{x}_{S})$ is the anchor point. The second term on the right-hand side corresponds to the value of the partial dependence function evaluated at the anchor point $\boldsymbol{x}_{S}^{a}$. Subtracting this constant value shifts the entire PDP so that it is zero at the anchor point, without altering its shape. The ICE curves are also anchored using the same principle. In the discussions and results that follow, we always depict the PDP and ICE curves anchored through this procedure. Both PDP and ICE may be misleading when dealing with correlated inputs. This work assumes that all inputs are independent, which is reasonable for design exploration or parameteric studies [23].